How many oxygen atoms are present in 4.90 mol in Al 203?


1
Expert's answer
2021-01-15T06:39:08-0500

Solution:

Al2O3 → 2Al + 3O

According to the above scheme: n(Al2O3) = n(Al)/2 = n(O)/3

n(O) = 3 × n(Al2O3) = 3 × 4.90 mol = 14.7 mol


One mole of any substance contains 6.022×1023 atoms/molecules.

Hence,

Number of O atoms = 14.7 mol × (6.022×1023 atoms / 1 mol) = 8.85×1024 O atoms

Number of O atoms = 8.85×1024 O atoms


Answer: 8.85×1024 oxygen atoms are present in 4.90 mol in Al2O3.